OpenStudy (anonymous):

Part A: Solve A = (x + 28) for x. (4 points) Part B: Determine the value of x when A = 135. (2 points) Part C: Solve -np - 40 > 10 for n. Show your work. (4 points)

OpenStudy (anonymous):

Part A: A-28=X A=(x+28) A-28=X because I isolated the 28 to the left side B: X=107 C: n<-40

OpenStudy (anonymous):

-np-40>10 -np/p>40 -n>40 n<-40 because you switch the signs when dividing by a negative number on inequalities.

OpenStudy (anonymous):

Thank You!